# Decision Quotient: A Regime-Sensitive Complexity Theory of Exact Relevance Certification

### 16 Mar 2026 v1

@@ 4.4 Stochastic Sufficiency
Definition 4.16 (Stochastic Anchor Sufficiency).
A coordinate set $I$ is stochastically anchor-sufficient if there exists an admissible fiber value $\alpha\in\mathrm{supp}(S_{I})$ and an action $a\in A$ such that
@@
Equivalently, some $I$-fiber has a unique conditional optimal action.
Problem 4.17 (STOCHASTIC-ANCHOR-SUFFICIENCY-CHECK).
@@
Question: Is $I$ stochastically anchor-sufficient?
Problem 4.18 (STOCHASTIC-MINIMUM-SUFFICIENT-SET).
@@
Question: Does there exist a stochastically decisive coordinate set $I$ with $|I|\leq k$?
4.5 Encoding-Sensitive Complexity of Stochastic Decisiveness
@@
Theorem 4.19 (Stochastic Decisiveness is Encoding-Sensitive).
For the stochastic decisiveness predicate above, Stochastic-Decisiveness-Check is decidable in polynomial time under the explicit-state encoding. Under the succinct encoding it is PP-hard. The accompanying upper-bound argument is proved in the paper via an explicit bad-fiber witness characterization whose finite witness/checking core is mechanized in the artifact.
(L: DC45, DC62, DC72, OU8-11.)
@@
For the upper bound, failure of decisiveness is witnessed by one bad state whose observed fiber does not have a singleton conditional optimum. The current artifact packages exactly this no-witness characterization together with the corresponding existential witness schema, and the two directions are bundled in a summary theorem (L: OU8-11.) These mechanized witness/checking packages verify the finite combinatorial core used by the paper’s oracle-class arguments. Concretely, the artifact verifies the existential-universal quantifier structure of the predicate, the witness-checking schema that guesses an anchor and verifies conditional-uniqueness using PP comparisons, and the finite-step-counted search wrappers (handles OU1 to OU2, OU6 to OU7) that bound the nondeterministic guessing phase. The standard oracle-machine reduction proving $\textsf{NP}^{\textsf{PP}}$ membership is argued in the paper text and not mechanized. The corresponding $\textsf{coNP}^{\textsf{PP}}$-style and $\textsf{NP}^{\textsf{PP}}$-style memberships are then proved in the paper text by standard complexity-theoretic reasoning.  $\blacksquare$
Proposition 4.20 (Stochastic Decisiveness Yields Stochastic Anchor Sufficiency).
If $I$ is stochastically decisive, then $I$ is stochastically anchor-sufficient.
(L: DC19.)
@@
$\text{Opt}^{\mathrm{stoch}}_{I}(\alpha)=\{a\}.$
@@
Theorem 4.21 (Stochastic Anchor Sufficiency is PP-hard).
Stochastic-Anchor-Sufficiency-Check is PP-hard.
(L: DC40-42.)
@@
Theorem 4.22 (Stochastic Minimum-Sufficient-Set is PP-hard).
Stochastic-Minimum-Sufficient-Set is PP-hard.
(L: DC45, DC47.)
@@
Proposition 4.23 (Explicit Finite Search for Stochastic Queries).
@@
Theorem 4.24 (Stochastic Anchor and Minimum Queries Lie in $\textsf{NP}^{\textsf{PP}}$).
@@
$\mathrm{STOCHASTIC\mbox{-}ANCHOR\mbox{-}SUFFICIENCY\mbox{-}CHECK}(P,I)\iff\exists s_{0}\,\exists a\;\bigl(\mathrm{fiberOpt}(P,I,s_{0})=\{a\}\bigr)$
is already available in the artifact PA3. A nondeterministic polynomial-time machine may therefore guess $(s_{0},a)$ and use a PP oracle to verify that $a$ is the unique conditional optimum on the anchor fiber. Concretely, uniqueness can be checked by comparing the conditional expected utility of $a$ against every competing action on that fiber, which is exactly the same majority-style conditional comparison used in the decisiveness verification.
For the minimum query, the machine guesses a coordinate set $I$ with $|I|\leq k$ and then asks whether $I$ is stochastically decisive. The same witness/checking template used for the anchor case yields the required PP-style verifier for the guessed set, so the minimum query is also in $\textsf{NP}^{\textsf{PP}}$.  $\blacksquare$
Theorem 4.24 completes the complexity picture for stochastic decisiveness: decisiveness is polynomial-time in the explicit-state model and PP-hard in the succinct model, while the anchor and minimum decisiveness queries are PP-hard and lie in $\textsf{NP}^{\textsf{PP}}$. Those oracle-class memberships are proved in the paper text. The artifact independently verifies the finite core behind those proofs: the bad-fiber counter-witness schema for decisiveness, the anchor-collapse equivalence, the generic existential-witness schema, bounded-witness recoveries for the explicit-state stochastic anchor and minimum searches, and existential-majority hardness/completeness packages for the anchor and decisiveness query families (L: OU8-11, EH2-10.)
@@
For stochastic preservation, this paper establishes explicit-state tractability for the preservation family, full-support inheritance, quotient equivalence, and support-sensitive obstruction/bridge lemmas beyond full support. The general-distribution minimum and anchor variants still remain open as full classification problems and are not claimed. For stochastic decisiveness, the paper establishes PP-hardness and $\textsf{NP}^{\textsf{PP}}$ membership under the stated encoding. The open entries in Table 2 are explicit scope boundaries rather than unresolved subcases.
The remaining gap is not a bookkeeping omission but a real structural issue. Stochastic decisiveness is a pure verification problem: once $I$ is fixed, one must check a family of majority-style conditional comparisons, which yields the stated encoding-sensitive complexity. By contrast, stochastic anchor and minimum sufficiency add an outer existential choice, namely to guess an anchor witness or a candidate coordinate set, before performing the same conditional verification. The current collapse theorem PA3 shows that anchor sufficiency reduces to existence of a fiber with a singleton conditional optimum, and the new existential-majority reduction shows that this extra witness layer is not vacuous. So the open question is genuinely whether that existential layer can be absorbed into PP under the present encoding, not whether the paper has simply stopped one step short of an obvious closure argument.
For that reason, the paper’s claim is intentionally calibrated. It does not present the stochastic row as one undifferentiated theorem about one predicate. Instead, it treats preservation and decisiveness as parallel exact questions. Preservation gives the direct stochastic analogue of the static question and the bridge back to the optimizer quotient. Decisiveness yields the strongest succinct-encoding complexity classification established in the paper. The split is part of the proved theory, not a temporary patch.
@@
The stochastic decisiveness query is fiber-indexed: fixing a candidate information set $I$ induces a derived decision problem whose optimizer is the conditional fiber optimizer. In the current formalization this induced problem is always $I$-sufficient (L: DC57), but its relevance notion is therefore also indexed by $I$. For that reason, the paper does not assert a single global relevance-set characterization for stochastic minimum sufficiency analogous to the static and sequential regimes.
4.6 Tractable Subcases
Proposition 4.25 (Tractable Stochastic Cases).
@@ 4.6 Tractable Subcases
4.7 Bridge from Static
@@
Proposition 4.26 (Static Sufficiency Does Not Imply Stochastic Sufficiency).
There exist instances where $I$ is sufficient in the static sense but not stochastically sufficient once a distribution $P$ is fixed.
(L: DQ58.)
@@
Choose utilities where optimizer ties are harmless pointwise but are broken in expectation after conditioning. Then static fibers preserve Opt, while the conditional optimizer obtained from $I$ fails to match the fully conditioned optimizer.  $\blacksquare$
Proposition 4.27 (Static Singleton Sufficiency Transfers to Stochastic Sufficiency).
Suppose that on every $I$-fiber the static optimal-action set is a singleton, and that this singleton is constant on that fiber. Then $I$ is stochastically sufficient for every distribution $P$.
(L: DQ15-16.)
@@
$\text{Opt}^{\mathrm{stoch}}_{I}(\alpha)=\{a_{\alpha}\}.$
Since this holds for every admissible fiber, the conditional optimizer on each fiber agrees with the fully conditioned optimizer at every state in that fiber. Hence $I$ is stochastically sufficient.  $\blacksquare$
The stochastic regime therefore gives the middle layer of the paper’s hierarchy: conditioning already raises exact certification beyond the static setting by turning exact certification into a counting comparison problem. Adding time will lift us further still: the next section shows how temporal contingency presents, alongside stochastic sufficiency, a decisiveness complexity family and then pushes the same underlying question from PP-style comparisons to PSPACE-level temporal certification.
5 Sequential Regime Complexity
@@
5.1 Sequential Decision Problems
Definition 5.1 (Sequential Decision Problem).
@@ 5.1 Sequential Decision Problems
Definition 5.2 (Sequential Sufficiency).
For a sequential instance $\mathcal{D}_{\mathrm{seq}}$, let $\text{Opt}_{\mathrm{seq}}(s)$ denote the optimal-action set of its induced decision map on state $s$. A coordinate set $I$ is sequentially sufficient if
$\forall s,s^{\prime}\in S:\quad s_{I}=s^{\prime}_{I}\implies\text{Opt}_{\mathrm{seq}}(s)=\text{Opt}_{\mathrm{seq}}(s^{\prime}).$
@@
Problem 5.3 (SEQUENTIAL-SUFFICIENCY-CHECK).
@@
Question: Is $I$ sequentially sufficient?
Definition 5.4 (Sequential Anchor Sufficiency).
A coordinate set $I$ is sequentially anchor-sufficient if there exists an anchor state whose $I$-agreement class preserves the optimal action set throughout that class.
Problem 5.5 (SEQUENTIAL-ANCHOR-SUFFICIENCY-CHECK).
@@
Question: Is $I$ sequentially anchor-sufficient?
Problem 5.6 (Sequential Minimum Query).
@@
Question: Does there exist a sequentially sufficient coordinate set $I$ with $|I|\leq k$?
Proposition 5.7 (Sequential Minimality Equals Relevance).
For any minimal sequentially sufficient set $I$, a coordinate lies in $I$ if and only if it is relevant for the underlying decision boundary.
(L: DC49-50.)
@@ 5.2 PSPACE-Completeness
Theorem 5.8 (Sequential Sufficiency is PSPACE-complete).
@@
Proposition 5.9 (Sequential Sufficiency Refines to Sequential Anchor Sufficiency).
If $I$ is sequentially sufficient, then $I$ is sequentially anchor-sufficient.
(L: DC23.)
@@
If optimal-action sets are preserved for every pair of states agreeing on $I$, then fixing any anchor state immediately yields preservation across its entire $I$-agreement class.  $\blacksquare$
Theorem 5.10 (Sequential Anchor Query is PSPACE-complete).
@@
Membership is in PSPACE. By definition, the query asks whether there exists an anchor state $s_{0}$ such that every state agreeing with $s_{0}$ on $I$ preserves the same optimal-action set. A nondeterministic polynomial-space procedure can guess $s_{0}$ and then scan candidate states one at a time, verifying the agreement condition and comparing the corresponding optimal-action sets. Since NPSPACE = PSPACE, the query lies in PSPACE.
@@
$\mathrm{TQBF}(q)\iff\mathrm{SEQUENTIAL\mbox{-}ANCHOR\mbox{-}SUFFICIENCY\mbox{-}CHECK}(\mathrm{reduceTQBF}(q),\emptyset).$
@@
Theorem 5.11 (Sequential Minimum Query is PSPACE-complete).
@@
Membership is in PSPACE. The query asks whether there exists a coordinate set $I$ of size at most $k$ such that $I$ is sequentially sufficient. A nondeterministic polynomial-space procedure can guess $I$ and then run the same polynomial-space verifier used for sequential sufficiency. Again, NPSPACE = PSPACE.
@@
Proposition 5.12 (Explicit Finite Search for Sequential Queries).
@@
5.3 Tractable Subcases
Proposition 5.13 (Tractable Sequential Cases).

### 31 Mar 2026 v2

@@ IV-D Stochastic Sufficiency
Definition IV.18 (Stochastic Anchor Sufficiency).
A coordinate set II is stochastically anchor-sufficient if there exists an admissible fiber value $\alpha\in\mathrm{supp}(S_{I})$ and an action $a\in A$ such that
@@
Equivalently, some II-fiber has a unique conditional optimal action.
Problem IV.19 (STOCHASTIC-ANCHOR-SUFFICIENCY-CHECK).
@@
Question: Is II stochastically anchor-sufficient?
Problem IV.20 (STOCHASTIC-MINIMUM-SUFFICIENT-SET).
@@
Question: Does there exist a stochastically decisive coordinate set II with $|I|\leq k$?
IV-E Encoding-Sensitive Complexity of Stochastic Decisiveness
@@
Theorem IV.21footnotemark: footnotetext: : DC45, DC62, DC72, OU8-11 (Stochastic Decisiveness is Encoding-Sensitive).
For the stochastic decisiveness predicate above, Stochastic-Decisiveness-Check is decidable in polynomial time under the explicit-state encoding. Under the succinct encoding it is $P$-hard. The accompanying upper-bound argument is proved in the paper via an explicit bad-fiber witness characterization whose finite witness/checking core is mechanized in the artifact.
@@
For the upper bound, failure of decisiveness is witnessed by one bad state whose observed fiber does not have a singleton conditional optimum. The current artifact packages exactly this no-witness characterization together with the corresponding existential witness schema, and the two directions are bundled in a summary theorem 262626: OU8-11. These mechanized witness/checking packages verify the finite combinatorial core used by the paper’s oracle-class arguments. Concretely, the artifact verifies the existential-universal quantifier structure of the predicate, the witness-checking schema that guesses an anchor and verifies conditional-uniqueness using $P$ comparisons, and the finite-step-counted search wrappers that bound the nondeterministic guessing phase.272727: OU1-2, OU6-7 The standard oracle-machine reduction proving $\textsf{NP}^{\textsf{PP}}$ membership is argued in the paper text and not mechanized. The corresponding $\textsf{coNP}^{\textsf{PP}}$-style and $\textsf{NP}^{\textsf{PP}}$-style memberships are then proved in the paper text by standard complexity-theoretic reasoning.  $\blacksquare$
Proposition IV.22footnotemark: footnotetext: : DC19 (Stochastic Decisiveness Yields Stochastic Anchor Sufficiency).
If II is stochastically decisive, then II is stochastically anchor-sufficient.
@@
$\mathrm{Opt}^{\mathrm{stoch}}_{I}(\alpha)=\{a\}.$
@@
Theorem IV.23footnotemark: footnotetext: : DC40-42 (Stochastic Anchor Sufficiency is $P$-hard).
Stochastic-Anchor-Sufficiency-Check is $P$-hard.
@@
Theorem IV.24footnotemark: footnotetext: : DC45, DC47 (Stochastic Minimum-Sufficient-Set is $P$-hard).
Stochastic-Minimum-Sufficient-Set is $P$-hard.
@@
Proposition IV.25footnotemark: footnotetext: : DC62-65, DC72-73, DC76, DC91-93 (Explicit Finite Search for Stochastic Queries).
@@
Theorem IV.26footnotemark: footnotetext: : OU1-2, OU6-7, PA3, EH1-5 (Stochastic Anchor and Minimum Queries Lie in $\textsf{NP}^{\textsf{PP}}$).
@@
$\mathrm{STOCHASTIC\mbox{-}ANCHOR\mbox{-}SUFFICIENCY\mbox{-}CHECK}(P,I)$
$\qquad\iff\exists s_{0}\,\exists a\;\bigl(\mathrm{fiberOpt}(P,I,s_{0})=\{a\}\bigr)$
is already available in the artifact. 333333: PA3 A nondeterministic polynomial-time machine may therefore guess $(s_{0},a)$ and use a $P$ oracle to verify that $a$ is the unique conditional optimum on the anchor fiber. Concretely, uniqueness can be checked by comparing the conditional expected utility of $a$ against every competing action on that fiber, which is exactly the same majority-style conditional comparison used in the decisiveness verification.
For the minimum query, the machine guesses a coordinate set II with $|I|\leq k$ and then asks whether II is stochastically decisive. The same witness/checking template used for the anchor case yields the required $P$-style verifier for the guessed set, so the minimum query is also in $\textsf{NP}^{\textsf{PP}}$.  $\blacksquare$
Theorem IV.26 completes the complexity picture for stochastic decisiveness: decisiveness is polynomial-time in the explicit-state model and $P$-hard in the succinct model, while the anchor and minimum decisiveness queries are $P$-hard and lie in $\textsf{NP}^{\textsf{PP}}$. Those oracle-class memberships are proved in the paper text. The artifact independently verifies the finite core behind those proofs: the bad-fiber counter-witness schema for decisiveness, the anchor-collapse equivalence, the generic existential-witness schema, bounded-witness recoveries for the explicit-state stochastic anchor and minimum searches, and existential-majority hardness/completeness packages for the anchor and decisiveness query families 343434: OU8-11, EH2-10.
@@
For stochastic preservation, this paper establishes explicit-state tractability for the preservation family, full-support inheritance, quotient equivalence, and support-sensitive obstruction/bridge lemmas beyond full support. The general-distribution minimum and anchor variants still remain open as full classification problems and are not claimed. For stochastic decisiveness, the paper establishes $P$-hardness and $\textsf{NP}^{\textsf{PP}}$ membership under the stated encoding. The open entries in Table 2 are explicit scope boundaries rather than unresolved subcases.
The open entries in Table 2 are explicit scope boundaries. The gap between $P$-hardness and $\textsf{NP}^{\textsf{PP}}$ membership for anchor/minimum decisiveness is structural, not a bookkeeping omission; resolving it requires determining whether the outer existential layer can be absorbed into $P$.
@@
The stochastic decisiveness query is fiber-indexed: fixing a candidate information set II induces a derived decision problem whose optimizer is the conditional fiber optimizer. In the current formalization this induced problem is always II-sufficient 353535: DC57, but its relevance notion is therefore also indexed by II. For that reason, the paper does not assert a single global relevance-set characterization for stochastic minimum sufficiency analogous to the static and sequential regimes.
IV-F Tractable Subcases
Proposition IV.27footnotemark: footnotetext: : DQ13, DQ15 (Tractable Stochastic Cases).
@@ IV-F Tractable Subcases
IV-G Bridge from Static
@@
Proposition IV.28footnotemark: footnotetext: : DQ58 (Static Sufficiency Does Not Imply Stochastic Sufficiency).
There exist instances where II is sufficient in the static sense but not stochastically sufficient once a distribution $P$ is fixed.
@@
Choose utilities where optimizer ties are harmless pointwise but are broken in expectation after conditioning. Then static fibers preserve $\mathrm{Opt}$, while the conditional optimizer obtained from II fails to match the fully conditioned optimizer.  $\blacksquare$
Proposition IV.29footnotemark: footnotetext: : DQ15-16 (Static Singleton Sufficiency Transfers to Stochastic Sufficiency).
Suppose that on every II-fiber the static optimal-action set is a singleton, and that this singleton is constant on that fiber. Then II is stochastically sufficient for every distribution $P$.
@@
$\mathrm{Opt}^{\mathrm{stoch}}_{I}(\alpha)=\{a_{\alpha}\}.$
Since this holds for every admissible fiber, the conditional optimizer on each fiber agrees with the fully conditioned optimizer at every state in that fiber. Hence II is stochastically sufficient.  $\blacksquare$
The stochastic regime therefore gives the middle layer of the paper’s hierarchy: conditioning already raises exact certification beyond the static setting by turning exact certification into a counting comparison problem. Adding time will lift us further still: the next section shows how temporal contingency presents, alongside stochastic sufficiency, a decisiveness complexity family and then pushes the same underlying question from $P$-style comparisons to PSPACE-level temporal certification.
V Sequential Regime Complexity
@@
V-A Sequential Decision Problems
Definition V.1 (Sequential Decision Problem).
@@ V-A Sequential Decision Problems
Definition V.2 (Sequential Sufficiency).
For a sequential instance $\mathcal{D}_{\mathrm{seq}}$, let $\mathrm{Opt}_{\mathrm{seq}}(s)$ denote the optimal-action set of its induced decision map on state $s$. A coordinate set II is sequentially sufficient if
$\forall s,s^{\prime}\in S:\quad s_{I}=s^{\prime}_{I}\implies\mathrm{Opt}_{\mathrm{seq}}(s)=\mathrm{Opt}_{\mathrm{seq}}(s^{\prime}).$
@@
Scope and research direction.
The sequential model studied here is a state-indexed product model, not a POMDP with belief-state dynamics or an MDP with policy optimization. This captures the case where each time step’s decision is governed by a time-indexed utility, and the question is which cross-temporal coordinates matter. Extending the optimizer quotient to belief-state POMDPs and history-dependent policies, where the quotient would be over belief states rather than world states and the complexity classification would need to engage with the policy optimization layer, remains a natural research direction.
Problem V.3 (SEQUENTIAL-SUFFICIENCY-CHECK).
@@
Question: Is II sequentially sufficient?
Definition V.4 (Sequential Anchor Sufficiency).
A coordinate set II is sequentially anchor-sufficient if there exists an anchor state whose II-agreement class preserves the optimal action set throughout that class.
Problem V.5 (SEQUENTIAL-ANCHOR-SUFFICIENCY-CHECK).
@@
Question: Is II sequentially anchor-sufficient?
Problem V.6 (Sequential Minimum Query).
@@
Question: Does there exist a sequentially sufficient coordinate set II with $|I|\leq k$?
Proposition V.7footnotemark: footnotetext: : DC49-50 (Sequential Minimality Equals Relevance).
For any minimal sequentially sufficient set II, a coordinate lies in II if and only if it is relevant for the underlying decision boundary.
@@ V-B PSPACE-Completeness
Theorem V.8footnotemark: footnotetext: : DC46, DC66, DC74 (Sequential Sufficiency is PSPACE-complete).
@@
Proposition V.9footnotemark: footnotetext: : DC23 (Sequential Sufficiency Refines to Sequential Anchor Sufficiency).
If II is sequentially sufficient, then II is sequentially anchor-sufficient.
@@
If optimal-action sets are preserved for every pair of states agreeing on II, then fixing any anchor state immediately yields preservation across its entire II-agreement class.  $\blacksquare$
Theorem V.10footnotemark: footnotetext: : DC28-29, DC44 (Sequential Anchor Query is PSPACE-complete).
@@
Membership is in PSPACE. By definition, the query asks whether there exists an anchor state $s_{0}$ such that every state agreeing with $s_{0}$ on II preserves the same optimal-action set. A nondeterministic polynomial-space procedure can guess $s_{0}$ and then scan candidate states one at a time, verifying the agreement condition and comparing the corresponding optimal-action sets. Since NPSPACE = PSPACE, the query lies in PSPACE.
@@
$\mathrm{TQBF}(q)$ $\iff$
$\mathrm{SEQUENTIAL\mbox{-}ANCHOR\mbox{-}SUFFICIENCY\mbox{-}CHECK}(\mathrm{reduceTQBF}(q),\emptyset).$
@@
Theorem V.11footnotemark: footnotetext: : DC46, DC48 (Sequential Minimum Query is PSPACE-complete).
@@
Membership is in PSPACE. The query asks whether there exists a coordinate set II of size at most $k$ such that II is sequentially sufficient. A nondeterministic polynomial-space procedure can guess II and then run the same polynomial-space verifier used for sequential sufficiency. Again, NPSPACE = PSPACE.
@@
Proposition V.12footnotemark: footnotetext: : DC66-69, DC74-76 (Explicit Finite Search for Sequential Queries).
@@
V-C Tractable Subcases
Proposition V.13footnotemark: footnotetext: : DQ12, DQ14, DQ45 (Tractable Sequential Cases).